OpenStudy (anonymous):

:work out these compliments .... what do you add to 2 and 1/8 to make 6 question is how do you work this out .. thx

OpenStudy (anonymous):

6 - 2 1/8 = 3 7/8

OpenStudy (anonymous):

2+(1/8) +x=6 solve for x

OpenStudy (anonymous):

let any number added to 2 and 1/8 to make 6 be y.then, 2+(1/8)+y=6 or,(17/8)+8y=48 or,y=5.734375